# Christoph Bandt

### v1

@@ 7 Construction of the topological space
Given an automaton $G,$ the definition of the space $X$ as quotient space does not provide much insight. We now present a meaningful construction of $X$ by approximation from the edge paths in $G.$ The approximating spaces look very much like graphs, as in Figure 5. They are finite topological spaces. A topology on a finite set is defined by assigning to each point $x$ a minimal open neighborhood $U_{x}$ [10, p. 2].
@@
Definition 2 (Approximations of automata-generated spaces)
Let $G$ be a topology-generating automaton with bounded equivalence classes and $X$ the generated topological space, as defined in Definition 1. Let $G_{k},k\in K$ be the automata of complete $k$-tuple addresses, as constructed in Section 6. For $k\in K$ and each level $n=1,2,...$ let $E^{n}_{k}$ denote the set of directed paths of $n$ edges in $G_{k}$ starting in $o.$ Each $y\in E_{k}^{n}$ corresponds to a set $W_{y}\subset D^{n}$ of $k$ words of length $n$ which form the labels of the path and are accepted at the endpoint of $y.$ Let
@@
We explain the notation and show that this is the natural definition. It is very similar to geometry of polyhedra with faces, edges, and vertices. The index set $K$ contains those $k$ for which complete $k$-tuples exist. For the square we had $K=\{2,4\},$ for Figures 4 and 5 we had $K=\{3\},$ for Example 3 we had $K=\{2,4,6,12\}.$ On the level $n,$ we consider the words $w$ of length $n,$ the corresponding disjoint cylinders $S_{w}$ in the symbolic space $S,$ and the pieces $X_{w}=\varphi(S_{w})$ of the topological space $X$ which are not disjoint. The pieces $X_{w}$ with $w\in D^{n}$ form a cover of $X.$ We turn this cover into a partition by taking all possible intersections of sets $X_{w}.$ These sets, the atoms of the partition, are considered as points of $X^{n}.$ They are given the topology which is inherited from the partition sets.
For the $2\times 2$ square, the $X_{w}$ are $4^{n}$ closed squares for $n=1,2,...$ Two neighboring squares will intersect in an edge, and four mutually neighboring squares have one common vertex point. The partition, shown in Figure 13, will consist of vertex points, edges without vertices, and open squares. The open squares correspond to $D^{n}$ and are open as points in $X^{n}.$ The minimal neighborhood $U_{x}$ of the edges contains the two neighboring open squares but no vertices. And the vertices are closed and their $U_{x}$ contains all four adjoining edges and four adjoining open squares.
For $n=1,2,...$ we get a sequence of finer and finer partitions with smaller and smaller sets which in the limit coincides with the space $X.$ The accurate concept is the inverse limit of the $X^{n}.$ There is a natural projection $\pi_{n+1,n}:X^{n+1}\to X^{n}$ which assigns to $(w_{1},...,w_{n+1})\in D^{n+1}$ the word $(w_{1},...,w_{n})$ in $D^{n},$ and to each path $y\in E_{k}^{n+1}$ the path of its first $n$ edges which belongs to $E_{k}^{n}.$ This projection is continuous, that is, for each open set in $X^{n}$ the corresponding refined set in $X^{n+1}$ is also open.
The inverse limit $X^{\infty}$ of $X^{1}{\pi_{2,1}\atop\leftarrow}X^{2}{\pi_{3,2}\atop\leftarrow}X^{3}...$ is the set of all sequences $x=(x^{1},x^{2},...)$ with $x^{n}\in X^{n}$ and $\pi_{n+1,n}(x^{n+1})=x^{n}.$ In our case the $x^{n}\in D^{n}$ extend to sequences $x\in D^{\infty}=S,$ and the edge paths $x^{n}\in E_{k}^{n}$ extend to infinite edge paths $x.$ Let $E_{k}^{\infty}$ denote the set of infinite edge paths in $G_{k}$ starting in $o,$ and
@@
Theorem 3 (Quotient space as inverse limit)
Let $G$ be a topology-generating automaton with bounded equivalence classes, $\varphi:S\to X$ the corresponding address map, and $X^{n}$ the finite spaces constructed above. Then $X$ is the compact Hausdorff space associated with the inverse limit $X^{\infty}.$
Proof. This statement mainly says that our construction of the approximations $X^{n}$ is mathematically correct. We explain how the elements of $E^{\infty}$ will disappear in the limit. A finite topological space can never be Hausdorff unless all sets are open. However, for an infinite space one usually requires that the intersection of all neighborhoods of some point $x$ is the point itself. In $X^{\infty}$ this holds for the points of $D^{\infty}$ but not for $x\in E^{\infty}.$ Such an $x$ represents an infinite edge path with $k$ address sequences $s_{1},...,s_{k}\in S\subset X^{\infty}.$ Each basic neighborhood of $x$ includes some finite part of these sequences in $X^{n}$ and thus includes $s_{1},...,s_{k}$ in $X^{\infty}.$ In order to obtain a Hausdorff space, we must identify $x$ with $s_{1},...,s_{k}.$ One way to define this identification is to say that $x\equiv y$ if $f(x)=f(y)$ for every continuous function on the space. This identification gets rid of $x$ and at the same time identifies the points $s_{1},...,s_{k}\in S.$ Thus we get exactly the quotient space $X$ as associated compact Hausdorff space of $X^{\infty}.$ To prove the homeomorphism, we state that all basic neighborhoods in $X^{\infty}$ obviously correspond to open sets in $X.$ On the other hand, given a neighborhood $V$ of $x\in X$ as a finite union of cylinder sets in $S,$ it is easy to determine an $n$ so that the minimal neighborhood $U(x^{n})$ leads to a subset of $V.$Box
Formally, the theorem remains true even if we would define the approximations $X^{n}$ only with the graph $G_{2},$ without multiple addresses. In that case the identification on $X^{\infty}$ would work because of the transitivity of the relation $\varphi(x)=\varphi(y).$ On finite stage, however, the identification could not be anticipated. For the square, for instance, we would have a hole in the center of $X^{1}$ and $(4^{n}-1)/3$ holes in $X^{n}$ but no holes in the limit.
@@ 8 Topological properties
The approximations $X_{n}$ are meaningful. The interval is characterized already by $X_{2}.$ Any open point is replaced by two open points and a closed centre point on the next level. Figure 13 shows a similar situation for the square. Each open point is replaced by $X_{1}$ in the next level, and each line segment spllits into two line segments with a middle point. The closed points do not split further.
It seems that topological properties of $X$ can be determined from $X_{n}$ on certain finite level $n.$ When we get into the cyclic part of $G,$ the changes of the $X_{n}$ will repeat those changes which were made for $X_{k}$ with $k<n.$ The structure of the approximations will not further change, it will only be refined. However, the critical $n$ depends on the topological property as well as on the automaton and can be difficult to determine. We briefly comment on some known results, open problems and work in progress.
@@
A classical result of Hata [22] says that $X$ is connected if and only if $X^{1}$ is connected. In the case of two digits, $X$ is connected if there is at least one double address [11]. A connected automata-generated space is locally connected and path connected, cf. [5, Section 3]. There seem to be no results on $k$-connectedness. A space is $k$-connected if for any two points $x,y$ there are $k$ connecting paths which are disjoint, except for the endpoints.
@@
For a connected $X,$ a point $x$ is a cutpoint if $X\setminus\{x\}$ is disconnected. If $x$ has a connected neighborhood $U,$ it is a local cutpoint if $U\setminus\{x\}$ is disconnected. Isolated loops in the automaton $G$ lead to local cutpoints, see Proposition 7. However, there are other cutpoints of a global nature, for instance the point with address $\overline{1}$ in Figure 5. The critical level for their detection could be $n=3.$ For tiles, the question was studied [29, 1] and for fractal squares in [41].
@@
There is a theory of topological dimension which can be applied also to finite spaces. Is it possible to determine the dimension of $X$ on finite level $n$? We can also ask for the smallest $n$ for which $X$ is embeddable in $\mathbb{R}^{n}.$ For graphs $H,$ there is a classical theorem of Kuratowski which says that $H$ is embeddable in the plane if and only if it has neither the complete graph $K_{5}$ nor the complete bipartite graph $K_{3,3}$ as homeomorphic subset. This fits well to finite spaces. We apply the simple part of this theorem to Example 2.
Proposition 8
Suppose some $X^{n}$ contains a discrete version of $K_{5}$ or $K_{3,3},$ where the edges are closed discrete arcs containing at least one point of $D^{n},$ and the vertices are projections of connected sets in $X.$ Then $X$ is not embeddable in plane. In particular, Example 2 is not embeddable in the plane.
Proof. An ordered set $x_{1},...,x_{n}$ is a discrete arc if for $k=1,...,n-1$ either $x_{k}$ is contained in the minimal neighborhood $U_{x_{k+1}}$ or $x_{k+1}$ is in $U_{x_{k}}.$ These are the images of $[0,1]$ under continuous mappings, and path connectedness in finite spaces is the same as connectedness [10]. If a graph $K$ is contained in $X^{n},$ we of course require that different arcs do not meet except at their endpoints. We made strong assumptions which imply that the copy of $K$ lifts to $X$ by the projection $\pi:X\to X^{n}.$ For each edge $e$ of $K,$ the set $\pi^{-1}(e)$ is a connected compact subset of $X$ which contains a piece $X_{w}$ of $X$ since $e$ contains a point of $D^{n}.$ For each vertex $c$ of $K,$ the set is $\pi^{-1}(c)$ is connected. We can choose a point from each $\pi^{-1}(c)$ and connect them by arcs in each $\pi^{-1}(e)$ to get a subset of $X$ homeomorphic to $K.$ Thus $X$ cannot be a subset of the plane.
For Example 2, $X^{n}$ contains $D^{n}$ and closed points $c$ which correspond to three equivalent addresses and join three open points. The open points are words $v$ in $D^{n}.$ A closed point will be denoted $vYw$ if it connects $v0w,v1w,v2w.$ According to the automaton, the possible cases are $vY,vY1,$ and $vY1w$ where the word $w$ contains only 0 and 2. Note that the closed points do not further split in $X^{m},m>n.$ They directly correspond to points in $X.$
We construct $K_{3,3}$ in $X^{4}.$ Let $012Y,112Y,212Y$ be the red vertices and $Y120,Y122$ and $110Y$ the blue vertices. We have to define an arc from each red vertex to each blue vertex. For the first two blue vertices, this arc consists of just one open point between the two endpoints. Namely, $012Y$ connects to $Y120$ via $0120,$ and to $Y122$ via $0122$ Similar for $112Y$ and $212Y.$ The connections to $110Y$ are a bit longer:
@@
Since each open point has only two adjacent closed points, and beside the vertices no point appears twice in our list, the edges are closed sets in $X^{4}$ which meet only at their endpoints. $\Box$
@@
In algebraic topology, finite spaces have shown to be a convenient tool [10, 17]. On the other hand, various attempts were made to determine algebraic invariants like the Euler characteristic for fractal spaces. However, the difference from a doughnut is that holes in fractals come in infinite families. So we must ask for characteristic exponents rather than absolute invariants. A first step would be to characterize contractible spaces which include trees [14] and disk-like tiles [1, 29, 31, 47].
@@ 9 Realizing automata as neighbor graphs of IFS
We now consider metric realizations of our spaces $X$ which already appeared in the Figures 3, 4, 7, 9, and 11. A finite set of one-to-one contractive maps $f_{0},...,f_{m-1}$ on $\mathbb{R}^{d}$ is called an iterated function system (IFS) [11]. There is a unique compact set $X\subset\mathbb{R}^{d}$ which fulfils $X=f_{0}(X)\cup...\cup f_{m-1}(X).$ This is the self-similarity equation (2) with prescribed homeomorphisms. Usually, the $f_{k}$ are assumed to be linear maps or even similitudes. As in [6, 7, 8] we assume that the $f_{k}$ are similitudes with the same similarity ratio $r<1.$ Compositions of the $f_{k}$ are expressed by words $u\in D^{n}$ as $f_{u}=f_{u_{1}}\cdot...\cdot f_{u_{n}}.$
@@
to the previously determined maps $h.$ We consider these maps as vertices of a graph and put an edge with label $(i,j)$ from $h$ to $h^{\prime}.$ The initial state is the identity map $id,$ and since $f_{j}^{-1}f_{j}=id,$ we have loops from $id$ to itself for every $j\in D.$
As described in Frougny and Sakarovich [18], there is a trim part of the graph, since by contractivity, maps $h$ with $|h(0)|>C$ for some constant $C$ will fulfil $|h^{\prime}(0)|>|h(0)|$ in (5) for all $i,j\in D.$ Such maps $h^{\prime}$ are neglected, as well as all vertices $h$ which lead only to such maps. If the maps are defined in a discrete setting, for instance a canonical number system [18] or an algebraic number field generated by a Pisot number [8], the trim part of the automaton will be finite, and the resulting automaton will fulfil the conditions of definition 1. In this case we shall say that the given IFS $f_{0},...,f_{m-1}$ is a linear or similitude representation of the automaton $G.$
@@
We provide no clear answers, but some methodology and examples. Let us start with the basic Example 1. The state right in Figure 1 is now called $h=f_{0}^{-1}f_{1},$ since the edge with label $(0,1)$ from $o=id$ leads to this state. We are looking for the unknown mappings $f_{0},f_{1}.$ The loop with label $(1,0)$ at this state now turns into an equation $h^{\prime}=h$ in (5):
$f_{1}^{-1}hf_{0}=h\ ,\mbox{ or }\ hf_{0}=f_{1}h\ .$ (6)
Such commutativity relations are typical conditions in our representation problem. We now change notation. Let $g$ be a expanding linear map with factor $R>1,$ and let $f_{k}=g^{-1}h_{k}$ for $k\in D.$ Then $f_{k}^{-1}=h_{k}^{-1}g,$ and $h_{k}$ is an isometry. Taking $g=f_{0}^{-1}$ and shifting the origin of our coordinate system we can assume that $g$ is a linear map and $h_{0}=id$ (cf. [8, Section 1]). Moreover, $f_{i}^{-1}f_{j}=h_{i}^{-1}gg^{-1}h_{j}=h_{i}^{-1}h_{j}.$
@@
The case of Figure 2b is a bit different. At state $b,$ we have $h_{1}=h_{1}^{-1}$ which implies $h_{1}(z)=-z+1$ where the number $b$ was standardized. Now $c=f_{1}^{-1}h_{1}f_{1}=h_{1}gh_{1}g^{-1}h_{1}$ and $c(z)=-z+2-\lambda.$ The loop at state $c$ reads $gcg^{-1}=c$ which leads to the quadratic equation $\lambda^{2}-3\lambda+2=0$ and $\lambda=2,c(z)=-z.$ For $X=[0,\frac{1}{2}],$ the two neighbor maps $h_{1}$ and $c$ are the point reflections at the endpoints of the interval. We proved a supplement to Proposition 4.
Proposition 9
For the automata in Figures 1 and 2, the only representations by complex linear maps are the IFS which belong to the number systems with base 2 and -2, and to the tent map. normal-$\Box$
The representation remained unique although we went from dimension one to two. However, the uniqueness is lost if we allow for orientation-reversing maps $g(z)=\lambda\overline{z}.$ Beside the binary interval, we then get Koch curves generated by two reflective maps.
For Example 3 it will now be shown that the whole algebra of the reflection group as well as the action of $g$ is contained in the automaton in Figure 11.
Proposition 10
@@
$a^{2}=b^{2}=c^{2}=id\ ,\ ab=ba\ ,\ (ac)^{3}=id\ ,\ (cb)^{6}=id\quad\mbox{ and }$
$gcg^{-1}=b\ ,\ gag^{-1}=cbc\ ,\ gbg^{-1}=cac=aca\ .$
Proof. We work in an abstract setting not restricted to $\mathbb{C}.$ It is crucial to set $h_{1}=id$ because $h_{0}=id$ would not lead to an orientation-preserving $g.$ The edges from $o=id$ then imply $c=h_{0}=h_{0}^{-1}$ and $a=h_{2}=h_{2}^{-1}.$ This shows $a^{2}=c^{2}=id.$ Now consider the two edges from $c$ to $b$ and the basic recursion (5) where $(i,j)$ are the edge labels.
$b=f_{1}^{-1}cf_{1}=gcg^{-1}\ ,\ b=f_{2}^{-1}cf_{2}=h_{2}gcg^{-1}h_{2}=aba\ .$
The first relation and $c^{2}=id$ yields $b^{2}=gcg^{-1}gcg^{-1}=id.$ The second relation gives $ab=ba,$ which is also self-inverse, so that we could conclude $(ab)^{2}=id.$ The edge from $a$ to $b$ is evaluated as
$b=f_{0}^{-1}af_{0}=cgag^{-1}c\ ,\ \mbox{ or }\ cbc=gag^{-1}\ .$
@@
Taking the third power, we obtain $(cb)^{6}=id.$ It is now clear that $ab,ac,$ and $bc$ are rotations and $a,b,c$ must be reflections. The first line in the proposition shows the well-known Coxeter relations for the reflection group generated by $a,b,c.$ The second line expresses the action of $g$ as an automorphism on that group, similar to a substitution. They do determine $g.$ If $X$ is the triangle formed by the reflection lines of $a,b,c,$ then the reflection lines of $cbc,cac,$ and $b$ determine another triangle which is $g(X)=a(X)\cup X\cup c(X).$Box
Conjecture. For any self-similar crystallographic or self-affine tile, the combinatorial structure of the neighbor automaton uniquely determines the IFS.

### v2

@@ 7 Construction of the topological space
Given an automaton $G$, the definition of the space $X$ as quotient space does not provide much insight. We now present a meaningful construction of $X$ by approximation from the edge paths in $G$. The approximating spaces look very much like graphs, as in Figure 5. They are finite topological spaces. A topology on a finite set is defined by assigning to each point $x$ a minimal open neighborhood $U_{x}$ [11, p. 2].
@@
Definition 7.1 (Approximations of automata-generated spaces).
Let $G$ denote a topology-generating automaton with finite equivalence classes, with size smaller than a constant $C$, and $X$ the generated topological space. Let $G_{k},k\in K$, be the automata of complete $k$-tuple addresses, as constructed in Section 6. For $k\in K$ and each level $n=1,2,\dots$ let $E^{n}_{k}$ denote the set of directed paths of $n$ edges in $G_{k}$ starting in $o$. Each $y\in E_{k}^{n}$ corresponds to a set $W_{y}\subset D^{n}$ of $k$ words of length $n$ which form the labels of the path and are accepted at the endpoint of $y$. Let
@@
We explain the notation and show that this is the natural definition. It is very similar to the geometry of polyhedra with faces, edges, and vertices. The index set $K$ contains those $k$ for which complete $k$-tuples exist. For the square we had $K=\{2,4\}$, for Figures 4 and 5 we had $K=\{3\}$, for Example 6.3 we had $K=\{2,4,6,12\}$. On the level $n$, we consider the words $w$ of length $n$, the corresponding disjoint cylinders $S_{w}$ in the symbolic space $S$, and the pieces $X_{w}=\varphi(S_{w})$ of the topological space $X$ which are not disjoint. The pieces $X_{w}$ with $w\in D^{n}$ form a cover of $X$. We turn this cover into a partition by taking all possible intersections of sets $X_{w}$. These sets, the atoms of the partition, are considered as points of $X^{n}$. They are given by the topology which is inherited from the partition sets.
For the $2\times 2$ square, the $X_{w}$ are $4^{n}$ closed squares for $n=1,2,\dots$. Two neighboring squares will intersect in an edge, and four mutually neighboring squares have one common vertex point. The partition, shown in Figure 13, will consist of vertex points, edges without vertices, and open squares. The open squares correspond to $D^{n}$ and are open as points in $X^{n}$. The minimal neighborhood $U_{x}$ of the edges contains the two neighboring open squares but no vertices. And the vertices are closed and their $U_{x}$ contains all four adjoining edges and four adjoining open squares.
For $n=1,2,\dots$. we get a sequence of finer and finer partitions with smaller and smaller sets which in the limit coincides with the space $X$. The accurate concept is the inverse limit of the $X^{n}$. There is a natural projection $\pi_{n+1,n}:X^{n+1}\to X^{n}$ which assigns to $(w_{1},\dots,w_{n+1})\in D^{n+1}$ the word $(w_{1},\dots,w_{n})$ in $D^{n}$, and to each path $y\in E_{k}^{n+1}$ the path of its first $n$ edges which belongs to $E_{k}^{n}$. This projection is continuous, that is, for each open set in $X^{n}$ the corresponding refined set in $X^{n+1}$ is also open.
[…]
$x=(x^{1},x^{2},\dots)\text{ with }x^{n}\in X^{n}\text{ and }\pi_{n+1,n}(x^{n+1})=x^{n}.$
In our case, the $x^{n}\in D^{n}$ extend to sequences $x\in D^{\infty}=S$, and the edge paths $x^{n}\in E_{k}^{n}$ extend to infinite edge paths $x$. Let $E_{k}^{\infty}$ denote the set of infinite edge paths in $G_{k}$ starting in $o$, and
@@
Theorem 7.2 (Quotient space as inverse limit).
Let $G$ be a topology-generating automaton with finite equivalence classes, with size smaller than a constant $C$. Let $\varphi:S\to X$ denote the address map, and $X^{n}$ the finite spaces constructed above. Then $X$ is the compact Hausdorff space associated with the inverse limit $X^{\infty}$.
Proof 7.3.
This statement mainly says that our construction of the approximations $X^{n}$ is mathematically correct. We explain how the elements of $E^{\infty}$ will disappear in the limit. A finite topological space can never be Hausdorff unless all sets are open. However, for an infinite space one usually requires that the intersection of all neighborhoods of some point $x$ is the point itself. In $X^{\infty}$ this holds for the points of $D^{\infty}$ but not for $x\in E^{\infty}$. Such an $x$ represents an infinite edge path with $k$ address sequences $s_{1},\dots,s_{k}\in S\subset X^{\infty}$. Each basic neighborhood of $x$ includes some finite part of these sequences in $X^{n}$ and thus includes $s_{1},\dots,s_{k}$ in $X^{\infty}$. In order to obtain a Hausdorff space, we must identify $x$ with $s_{1},\dots,s_{k}$. One way to define this identification is to say that $x\equiv y$ if $f(x)=f(y)$ for every continuous function on the space. This identification gets rid of $x$ and at the same time identifies the points $s_{1},\dots,s_{k}\in S$. Thus we get exactly the quotient space $X$ as associated compact Hausdorff space of $X^{\infty}$. To prove the homeomorphism, we state that all basic neighborhoods in $X^{\infty}$ obviously correspond to open sets in $X$. On the other hand, given a neighborhood $V$ of $x\in X$ as a finite union of cylinder sets in $S$, it is easy to determine an $n$ so that the minimal neighborhood $U(x^{n})$ leads to a subset of $V$.
Formally, the theorem remains true even if we would define the approximations $X^{n}$ only with the graph $G_{2}$, without multiple addresses. In that case the identification on $X^{\infty}$ would work because of the transitivity of the relation $\varphi(x)=\varphi(y)$. On the finite stage, however, the identification could not be anticipated. For the square, for instance, we would have a hole in the center of $X^{1}$ and $(4^{n}-1)/3$ holes in $X^{n}$ but no holes in the limit.
@@ 8 Topological properties
The approximations $X^{n}$ are meaningful. The interval is characterized already by $X^{2}$. Any open point is replaced by two open points and a closed centre point on the next level. Figure 13 shows a similar situation for the square. Each open point is replaced by $X^{1}$ in the next level, and each line segment splits into two line segments with a middle point. The closed points do not split further.
It seems that topological properties of $X$ can be determined from $X^{n}$ on certain finite level $n$. When we get into the cyclic part of $G$, the changes of the $X^{n}$ will repeat those changes which were made for $X^{k}$ with $k<n$. The structure of the approximations will not further change, it will only be refined. However, the critical $n$ depends on the topological property as well as on the automaton and can be difficult to determine. We briefly comment on some known results, open problems and work in progress.
@@
A classical result of Hata [23] says that $X$ is connected if and only if $X^{1}$ is connected. In the case of two digits, $X$ is connected if there is at least one double address [12]. A connected automata-generated space is locally connected and path connected, cf. [7, Section 3]. There seem to be no results on $k$-connectedness. A space is $k$-connected if for any two points $x,y$ there are $k$ connecting paths which are disjoint, except for the endpoints.
@@
For a connected $X$, a point $x$ is a cutpoint if $X\setminus\{x\}$ is disconnected. If $x$ has a connected neighborhood $U$, it is a local cutpoint if $U\setminus\{x\}$ is disconnected. Isolated loops in the automaton $G$ lead to local cutpoints, see Proposition 5.1. However, there are other cutpoints of a global nature, for instance the point with address $\overline{1}$ in Figure 5. The critical level for their detection could be $n=3$. For tiles, the question was studied [29, 1] and for fractal squares in [43].
@@
There is a theory of topological dimension which can be applied also to finite spaces. Is it possible to determine the dimension of $X$ on finite level $n$? We can also ask for the smallest $n$ for which $X$ is embeddable in $\mathbb{R}^{n}$. For graphs $H$, there is a classical theorem of Kuratowski which says that $H$ is embeddable in the plane if and only if it has neither the complete graph $K_{5}$ nor the complete bipartite graph $K_{3,3}$ as homeomorphic subset. This fits well with finite spaces. We apply the simple part of this theorem to Example 3.4.
Proposition 8.1.
Suppose some $X^{n}$ contains a discrete version of $K_{5}$ or $K_{3,3}$, where the edges are closed discrete arcs containing at least one point of $D^{n}$, and the vertices are projections of connected sets in $X$. Then $X$ is not embeddable in the plane. In particular, Example 3.4 is not embeddable in the plane.
Proof 8.2.
An ordered set $x_{1},\dots,x_{n}$ is a discrete arc if for $k=1,\dots,n-1$ either $x_{k}$ is contained in the minimal neighborhood $U_{x_{k+1}}$ or $x_{k+1}$ is in $U_{x_{k}}$. These are the images of $[0,1]$ under continuous mappings, and path connectedness in finite spaces is the same as connectedness [11]. If a graph $K$ is contained with discrete arcs in $X^{n}$, we of course require that different arcs do not meet except at their endpoints. We made strong assumptions concerning the lift of $K$ to $X$ by the projection $\pi:X\to X^{n}$. For each edge $e$ of $K$, the set $\pi^{-1}(e)$ is a connected compact subset of $X$ which contains a piece $X_{w}$ of $X$ since $e$ contains a point of $D^{n}$. For each vertex $c$ of $K$, the set $\pi^{-1}(c)$ is connected. We can choose a point from each $\pi^{-1}(c)$ and connect them by arcs in each $\pi^{-1}(e)$ to get a subset of $X$ which topologically realizes the graph $K$ with proper arcs. Thus $X$ cannot be a subset of the plane.
In Example 3.4, the equivalence relation is given only by triple addresses, indicated as triangles in Figure 5. Three pieces of $X$ corresponding to words $v0w,v1w,v2w$ will intersect, according to Figure 10. Here $v$ can be any word, and $w$ can be empty, $1$, or $1u$ with $u\in\{0,2\}^{*}$. On level $n$, the pieces are represented by open points in $D^{n}$, and their intersection is represented as a closed point in $X^{n}$ which we denote $vYw$ since it is in the closure of $v0w$, of $v1w$, and of $v2w$. Note that the closed points do not further split in $X^{m},m>n$. They directly correspond to points in $X$.
We construct $K_{3,3}$ in $X^{4}$. Let $012Y,112Y,212Y$ be the red vertices (with empty $w$) and $Y120,Y122$ and $110Y$ the blue vertices. We have to define an arc from each red vertex to each blue vertex. For the first two blue vertices, this arc consists of just one open point between the two endpoints. Namely, $012Y$ connects to $Y120$ via $0120$, and to $Y122$ via $0122$. Similar for $112Y$ and $212Y$. The connections to $110Y$ are a bit longer:
@@
Since each open point has only two adjacent closed points, and beside the vertices no point appears twice in our list, the edges are closed sets in $X^{4}$ which meet only at their endpoints.
@@
In algebraic topology, finite spaces have become a convenient tool [11, 18]. On the other hand, various attempts were made to determine algebraic invariants like the Euler characteristic for fractal spaces. However, the difference from a doughnut is that holes in fractals come in infinite families. So we must ask for characteristic exponents rather than absolute invariants. A first step would be to characterize contractible spaces which include trees [15] and disk-like tiles [1, 29, 32, 49].
@@ 9 Realizing automata as neighbor graphs of IFS
We now consider metric realizations of our spaces $X$ which already appeared in the Figures 3479, and 11. A finite set of one-to-one contractive maps $f_{0},\dots,f_{m-1}$ on $\mathbb{R}^{d}$ is called an iterated function system (IFS) [12]. There is a unique compact set $X\subset\mathbb{R}^{d}$ which fulfils $X=f_{0}(X)\cup\dots\cup f_{m-1}(X)$. This is the self-similarity equation (2) with prescribed homeomorphisms. Usually, the $f_{k}$ are assumed to be linear maps or even similitudes. As in [8, 9, 6] we assume that the $f_{k}$ are similitudes with the same similarity ratio $r<1$. Compositions of the $f_{k}$ are expressed by words $u\in D^{n}$ as $f_{u}=f_{u_{1}}\cdot\dots\cdot f_{u_{n}}$.
@@
to the previously determined maps $h$. We consider these maps as vertices of a graph and put an edge with the label $(i,j)$ from $h$ to $h^{\prime}$. The initial state is the identity map $id$, and since $f_{j}^{-1}f_{j}=id$, we have loops from $id$ to itself for every $j\in D$.
As described in Frougny and Sakarovich [19], there is a trim part of the graph, since by contractivity, maps $h$ with $|h(0)|>C$ for some constant $C$ will fulfil $|h^{\prime}(0)|>|h(0)|$ in (5) for all $i,j\in D$. Such maps $h^{\prime}$ are neglected, as well as all vertices $h$ which lead only to such maps. If the maps are defined in a discrete setting, for instance a canonical number system [19] or an algebraic number field generated by a Pisot number [6], the trim part of the automaton will be finite, and the resulting automaton will fulfil the conditions of definition 2.4. In this case we shall say that the given IFS $f_{0},\dots,f_{m-1}$ is a linear or similitude representation of the automaton $G$.
@@
We provide no clear answers, but some methodology and examples. Let us start with the basic Example 2.1. The state right in Figure 1 is now called $h=f_{0}^{-1}f_{1}$, since the edge with label $(0,1)$ from $o=id$ leads to this state. We are looking for the unknown mappings $f_{0},f_{1}$. The loop with label $(1,0)$ at this state now turns into an equation $h^{\prime}=h$ in (5):
$f_{1}^{-1}hf_{0}=h\,\mbox{ or }\ hf_{0}=f_{1}h.$ (6)
Such commutativity relations are typical conditions in our representation problem. We now change the notation. Let $g$ be an expanding linear map with factor $R>1$, and let $f_{k}=g^{-1}h_{k}$ for $k\in D$. Then $f_{k}^{-1}=h_{k}^{-1}g$, and $h_{k}$ is an isometry. Taking $g=f_{0}^{-1}$ and shifting the origin of our coordinate system we can assume that $g$ is a linear map and $h_{0}=id$ (cf. [6, Section 1]). Moreover, $f_{i}^{-1}f_{j}=h_{i}^{-1}gg^{-1}h_{j}=h_{i}^{-1}h_{j}$.
@@
The case of Figure 2b is a bit different. At state $b$, we have $h_{1}=h_{1}^{-1}$ which implies $h_{1}(z)=-z+1$ where the number $b$ was standardized. Now $c=f_{1}^{-1}h_{1}f_{1}=h_{1}gh_{1}g^{-1}h_{1}$ and $c(z)=-z+2-\lambda$. The loop at state $c$ reads $gcg^{-1}=c$ which leads to the quadratic equation $\lambda^{2}-3\lambda+2=0$ and $\lambda=2,c(z)=-z$. For $X=[0,\frac{1}{2}]$, the two neighbor maps $h_{1}$ and $c$ are the point reflections at the endpoints of the interval. We proved a supplement to Proposition 3.1.
Proposition 9.1.
For the automata in Figures 1 and 2, the only representations by complex linear maps are the IFS which belong to the number systems with base 2 and -2, and to the tent map. $\Box$
The representation remained unique although we went from dimension one to two. However, the uniqueness is lost if we allow for orientation-reversing maps $g(z)=\lambda\overline{z}$. Besides the binary interval, we then get Koch curves generated by two reflective maps.
For Example 6.3 it will now be shown that the whole algebra of the reflection group as well as the action of $g$ is contained in the automaton in Figure 11.
Proposition 9.2.
@@
$a^{2}=b^{2}=c^{2}=id\,\ ab=ba\,\ (ac)^{3}=id\,\ (cb)^{6}=id\quad\mbox{ and }$
$gcg^{-1}=b\,\ gag^{-1}=cbc\,\ gbg^{-1}=cac=aca.$
Proof 9.3.
We work in an abstract setting not restricted to $\mathbb{C}$. It is crucial to set $h_{1}=id$ because $h_{0}=id$ would not lead to an orientation-preserving $g$. The edges from $o=id$ then imply $c=h_{0}=h_{0}^{-1}$ and $a=h_{2}=h_{2}^{-1}$. This shows $a^{2}=c^{2}=id$. Now consider the two edges from $c$ to $b$ and the basic recursion (5) where $(i,j)$ are the edge labels.
$b=f_{1}^{-1}cf_{1}=gcg^{-1}\,\ b=f_{2}^{-1}cf_{2}=h_{2}gcg^{-1}h_{2}=aba.$
The first relation and $c^{2}=id$ yields $b^{2}=gcg^{-1}gcg^{-1}=id$. The second relation gives $ab=ba$, which is also self-inverse so that we could conclude $(ab)^{2}=id$. The edge from $a$ to $b$ is evaluated as
$b=f_{0}^{-1}af_{0}=cgag^{-1}c\,\ \mbox{ or }\ cbc=gag^{-1}.$
@@
Taking the third power, we obtain $(cb)^{6}=id$. It is now clear that $ab,ac$, and $bc$ are rotations and $a,b,c$ must be reflections. The first line in the proposition shows the well-known Coxeter relations for the reflection group generated by $a,b,c$. The second line expresses the action of $g$ as an automorphism on that group, similar to a substitution. They do determine $g$. If $X$ is the triangle formed by the reflection lines of $a,b,c$, then the reflection lines of $cbc,cac$, and $b$ determine another triangle which is $g(X)=a(X)\cup X\cup c(X)$.
Conjecture 9.4.
For any self-similar crystallographic or self-affine tile, the combinatorial structure of the neighbor automaton uniquely determines the IFS.